Group Selection Page
HELP
Controls Group
Data Analysis Group
ITDM
WP76
DET
Vacuum
EEE
WP78
WP74
FXE
SCS
SPB
MID
HED
SQS
SXP
Sample Environment
Operation
PSPO
XO
TS
Migrated2Zulip
|
Technical integration and coordination
SASE1 Status and SASE3 Readiness
|
Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther |
Not logged in |
 |
|
Message ID: 570
Entry time: 02 Dec 2024, 09:56
|
Author: |
Frederik Wolff-Fabris |
Location: |
General |
Severity: |
Information |
Subject: |
Notes XO Integration Meeting 02.12.2024 |
|
|
General Announcement:
- A. Galler informs on the changed schedule for Operation due to the difficulties with the cry plant last week: the current beamtimes are extended until Wed 07:00 and then tuning for cw49 will take place.
It is possible to have special arrangements between instruments - in such case please let XO know.
Expectation is to have delivery from 15:00 on Wednesday to cw49 delivery, but depends on the promptness of the tuning.
- A. Silenzi announces that DOC has scheduled one extra day of support on Tuesday from 07:00 to 23:00.
CTRL:
- Nothing to report.
DA:
- Nothing to report.
EEE:
- Nothing to report.
VAC:
- Nothing to report.
XPD:
- Everything fine with Diagnostics.
DET:
- Nothing to report.
ITDM:
- Nothing to report.
XRO:
- Nothing to report.
LAS:
- OCD call from Desy looking for SA3 instrument to change repletion rate - to be clarified.
UNSYS:
- Cell20 SA3 phase shifter motor has a problem. Repair postponed to the winter maintenance period.
SPB:
- Using the beam until Wednesday morning.
FXE:
- Nothing to report.
MID:
- All fine.
HED:
- Not present.
SCS:
- Nothing to report.
SQS:
- 2CPP experiment: unexpected changes happened on Saturday when SA2 was tuned.
- R. Villanueva as PRC comments that a tuning to recover SA1 intensity is under discussion now and being coordinated.
SXP:
- Nothing to report. |